# Adaptive Plan-Execute Framework for Smart Contract Security Auditing

### v1

@@ Abstract
Large Language Models (LLMs) have shown great promise in code analysis and auditing; however, they still struggle with hallucinations and limited context-aware reasoning. We introduce SmartAuditFlow, a novel Plan-Execute framework that enhances smart contract security analysis through dynamic audit planning and structured execution. Unlike conventional LLM-based auditing approaches that follow fixed workflow, predefined steps, SmartAuditFlow dynamically generates and refines audit plans based on the unique characteristics of each smart contract. It continuously adjusts its auditing strategy in response to intermediate LLM outputs and newly detected vulnerabilities, ensuring a more adaptive and precise security assessment. The framework then executes these plans step by step, applying a structured reasoning process to enhance vulnerability detection accuracy while minimizing hallucinations and false positives. To further improve audit precision, SmartAuditFlow integrates iterative prompt optimization and external knowledge sources, such as static analysis tools and Retrieval-Augmented Generation (RAG). This ensures audit decisions are contextually informed and backed by real-world security knowledge, producing comprehensive security reports. Extensive evaluations across multiple benchmarks demonstrate that SmartAuditFlow outperforms existing methods, achieving 100% accuracy on common and critical vulnerabilities, 41.2% accuracy for comprehensive coverage of known smart contract weaknesses in real-world projects, and successfully identifying all 13 tested CVEs. These results highlight SmartAuditFlow’s scalability, cost-effectiveness, and superior adaptability over traditional static analysis tools and contemporary LLM-based approaches, establishing it as a robust solution for automated smart contract auditing.
@@ 1 Introduction
Existing smart contract security auditing primarily relies on manual code reviews by security experts, often supplemented by automated analysis tools zhang2023demystifying. While valuable, these approaches face significant limitations in the fast-paced blockchain environment. Manual audits, despite their potential for depth, are inherently time-consuming and costly, often creating bottlenecks in development cycles. Furthermore, the continuous evolution of blockchain technology and the emergence of novel attack vectors challenge the ability of human auditors to consistently identify all vulnerabilities sun2024gptscan. Expert reviews also rely on individual experience and knowledge, which can introduce subjectivity and gaps in coverage. Conversely, conventional automated tools, while offering speed, frequently suffer from high false positive rates and struggle to detect complex, context-dependent vulnerabilities. Such tools often lack the nuanced understanding of business logic and intricate blockchain-specific security patterns required to identify subtle flaws, such as sophisticated logic errors or vulnerabilities arising from inter-contract interactions, thereby limiting their overall effectiveness.
@@
The advent of Large Language Models (LLMs) presents a promising avenue for advancing smart contract auditing and addressing the aforementioned limitations liu2024propertygpt. LLMs possess the unique ability to process and understand both natural language (e.g., contract specifications, comments) and programming languages (i.e., smart contract code) liu2023your. This facilitates a deeper reasoning about the alignment between a contract’s intended functionality and its actual implementation. Their notable capabilities in pattern recognition and contextual analysis are particularly valuable for detecting subtle or novel security vulnerabilities that traditional automated tools might overlook. Furthermore, LLMs are good at reasoning processes and explaining identified security issues in an accessible manner, thereby empowering developers who may not be security specialists to understand and address the findings.
@@
Recent research has shown encouraging results in applying LLMs to smart contract security analysis david2023you; wei2023survey. Nevertheless, substantial challenges hinder the development of fully reliable and effective LLM-based auditing systems. Key issues include: (1) output inconsistency and non-determinism bender2021dangers; (2) the difficulty of validating the correctness and completeness of LLM-generated security assessments; and (3) limitations in ensuring comprehensive exploration of potential vulnerabilities. These factors necessitate careful strategies to harness LLM strengths while mitigating their inherent weaknesses. These challenges can be effectively addressed by adopting a workflow-based approach. This methodology decomposes the complex auditing task into manageable, discrete steps, facilitating structured and iterative interactions with LLMs. The efficacy of such structured methods is evident in the improved performance of advanced LLMs in complex reasoning when guided by systematic procedures zhong2024evaluation; guan2024deliberative, and is central to agentic workflows like HuggingGPT, Manus, and AutoGPT shen2023hugginggpt; wang2024survey, which orchestrate LLMs for sophisticated problem-solving.
@@ 2.1 Smart Contract Auditing
The rapid growth of smart contracts across blockchain networks has brought transformative capabilities, particularly to the Decentralized Finance (DeFi) sector. This rapid adoption, however, has simultaneously highlighted the critical importance of robust security measures, as vulnerabilities within smart contracts have frequently led to substantial financial losses tolmach2021survey; wei2023survey. The DeFi market, for instance, reached a peak total value locked (TVL) of approximately \$179 billion USD on November 9, 2021, with the Ethereum network alone contributing \$108.9 billion USD (around 60% of the total DeFi TVL at the time). As these blockchain-based applications attract significant capital and user engagement, they have also become prime targets for malicious actors. Until the end of 2024, the total amount of money lost by blockchain hackers exceeds 35.32 billion USD, stemming from more than 1800 hacks. A particularly severe recent incident occurred in February 2025, involving the theft of approximately \$1.5 billion USD in Ethereum from the Bybit digital asset exchange, reportedly one of the largest single cryptocurrency heists to date  [^2] https://www.csis.org/analysis/bybit-heist-and-future-us-crypto-regulation.
A smart contract audit is a meticulous and systematic security assessment designed to identify and mitigate vulnerabilities, logical errors, and inefficiencies within the contract’s codebase. Different from foundational blockchain security concerning consensus mechanisms or virtual machines, smart contract security operates at the application layer, which may be more irregular and complex. During an audit, security experts scrutinize all pertinent materials, including source code, whitepapers, and architectural documentation. This process typically involves manual inspection to uncover nuanced flaws that automated tools might miss, complemented by the use of specialized static and dynamic analysis tools to detect common vulnerability patterns and deviations from best practices ma2024combining; xia2024auditgpt. The cost of such audits can range from \$5,000 to over \$50,000 USD, contingent on contract complexity, and demand deep expertise in software engineering, blockchain-specific programming languages (like Solidity), and DeFi protocols. Audit durations can vary from a few days for simple contracts to several weeks or months for complex decentralized applications (dApps).
@@ 2.2 Automated Vulnerability Detection
Traditional automated techniques for smart contract vulnerability detection, including static analysis, symbolic execution, and fuzz testing, have been instrumental in early efforts to secure smart contracts chaliasos2024smart; durieux2020empirical. However, these methods face significant challenges related to scalability, computational efficiency, and their ability to comprehensively analyze increasingly complex smart contract systems. A critical study by Zhang et al. zhang2023demystifying revealed that 37 widely-used automated detection tools failed to identify over 80% of 516 exploitable vulnerabilities found in 167 real-world smart contracts, underscoring substantial blind spots in conventional approaches. The study also highlighted that the intricacies of detecting complex bugs often necessitate the collaborative efforts of multiple human auditors, further emphasizing the limitations of standalone automated tools and the high cost of thorough manual reviews.
Recent advancements in LLMs offer a promising frontier for enhancing smart contract security analysis david2023you; wei2023survey. LLMs demonstrate strong capabilities in understanding and reasoning about both natural language and programming code. This allows them to potentially bridge the gap between intended contract behavior and actual implementation logic, identifying vulnerabilities, verifying compliance with specifications, and assessing logical correctness with a degree of nuance previously unattainable by purely algorithmic tools. LLMs can excel at reasoning about complex security scenarios, explaining identified vulnerabilities in an understandable manner, evaluating business logic intricacies, and even hypothesizing novel attack vectors chen2023chatgpt; sun2024gptscan. Studies exploring LLMs in the smart contract domain have reported promising results, in some instances demonstrating performance comparable or superior to traditional automated tools. However, the probabilistic nature of LLMs means that their outputs can be inconsistent, and developing robust mechanisms for validating their findings remains an active area of research and a critical consideration for practical deployment azamfirei2023large; chen2023diversevul.
@@ 2.3 Agentic Systems and Workflows
Effectively harnessing LLMs for complex tasks like smart contract auditing requires sophisticated interaction strategies beyond simple querying. Prompt engineering, the art of crafting effective inputs to guide LLM outputs, is a cornerstone of LLM application. It is often more computationally efficient and cost-effective for adapting pre-trained LLMs to specific tasks than full model retraining or extensive fine-tuning, especially for rapid prototyping and iteration. However, LLMs are highly sensitive to prompt phrasing; minor variations in input can lead to significantly different, sometimes degraded, performance. For instance, one study demonstrated accuracy drops of up to 65% across several advanced LLMs due to subtle changes in question framing mirzadeh2024gsm. To address this sensitivity and enhance LLM reasoning, researchers have developed advanced prompting methodologies. These include techniques like chain-of-thought (CoT) prompting, which encourages step-by-step reasoning li2023camel; few-shot learning, which provides examples within the prompt wang2024chain; and more complex structured reasoning frameworks like Tree of Thought (ToT) yao2024tree and Graph of Thought (GoT) besta2024graph. These methods aim to elicit more reliable and accurate responses by structuring the LLM’s generation process.
Beyond one attempt advanced prompting, more sophisticated operational paradigms are emerging to tackle multi-step, complex problems. Two prominent approaches are autonomous agents and agentic workflows. Autonomous agents are designed to perceive their environment, make decisions, and take actions dynamically to achieve goals, often involving flexible, self-directed reasoning and tool use zhugegptswarm. They typically require careful design of their action space and decision-making heuristics tailored to their operational context. In contrast, agentic workflows structure problem-solving by decomposing a complex task into a sequence of predefined or dynamically chosen sub-tasks, often involving multiple LLM invocations, potentially with specialized roles or prompts for each step hong2024data; zhang2024aflow. This paradigm emphasizes a more explicit orchestration of LLM capabilities, leveraging human domain expertise in designing the workflow structure and allowing for iterative refinement of both the process and the intermediate results. Agentic workflows can integrate external tools, knowledge bases, and feedback loops, making them well-suited for tasks requiring methodical execution and high reliability, such as security auditing.
@@ 3.2 Plan Phase
The Plan phase is critical for preparing focused and effective tasks for the subsequent Execute Phase. Central to this is guiding the LLM using precisely engineered prompts. Prompts fundamentally shape the LLM’s interpretation of smart contract structures, identification of high-risk functions, and overall focus during analysis. As highlighted by prior research, poorly constructed or misaligned prompts can significantly degrade LLM output quality, potentially leading to incomplete audits or misinterpretation of vulnerabilities zhong2024evaluation; guan2024deliberative. Therefore, the systematic generation and selection of optimal prompts for each reasoning step is crucial.
@@ 3.2.1 Initial Analysis and Adaptive Planning Stages (A1-A2)
To achieve A1 and A2 stages, we employ an iterative prompt optimization strategy inspired by recent advancements in automatic prompt engineering weng2022large; zhou2022large; cheng2023black. This strategy aims to discover an optimal instruction $\rho^{*}$ that, when prepended to a given smart contract code $c$, maximizes the quality of the LLM’s output $A$ (e.g., a comprehensive initial analysis or a robust audit plan). Moreover, Stage A1 is enhanced by integrating insights from static analysis tools to ground the LLM’s assessments and improve the accuracy of its initial findings. These insights inform the subsequent prompt generation and planning in later stages.
@@ 3.2.2 Iterative Prompt Optimization Framework
  • Evolution Loop with Mini-Batch Optimization li2014efficient; stapor2022mini: Let $\mathcal{U}_{t}$ denote the instruction set at generation $t$. Each iteration of the loop ($t=1\dots T$) proceeds as follows:
@@
This component measures how well the LLM’s generated output $\mathcal{M}(\rho,c)$ aligns with the expected output $A$ from the dataset $\mathcal{D}_{task}$. For smart contract auditing tasks, $A$ might represent an expert-identified set of vulnerabilities, a correctly summarized contract behavior, or a well-structured audit plan component. Alignment can be measured using metrics like BERTScore, ROUGE, or BLEU vaswani2017attention; es2024ragas for textual similarity if $A$ is text, or more domain-specific metrics. We define a domain-specific accuracy metric, drawing upon best practices and guidelines from leading smart contract audit firms (e.g., ConsenSys Diligence, Trail of Bits, or OpenZeppelin).
@@ 3.2.3 Static Analysis Tool Involvement
Widely recognized tools such as Slither, Mythril, Solint, and Oyente automatically detect common vulnerability patterns. However, their utility in our framework extends beyond standalone detection. Specifically, within Stage A1, the outputs from these tools are leveraged to significantly enhance the LLM’s judgements. Among these tools, we selected slither-0.11.0 as our static analysis tool. This choice is motivated by Slither’s active maintenance, frequent updates driven by a large developer community, and its comprehensive suite of nearly 100 distinct vulnerability detectors. In contrast, tools like Mythril primarily focus on low-level EVM bytecode analysis, which, while valuable, offers a different granularity of insight, and other tools such as Solint and Oyente have seen limited updates in recent years. Slither’s proficiency in extracting detailed code syntax and semantic information (e.g., inheritance graphs, function call graphs, state machine representations) has been well-documented in prior research huang2023semantic; ayub2023storage.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in smart contract auditing. However, they are still susceptible to hallucinations and limited context-aware reasoning. In this paper, we propose SmartAuditFlow, a dynamic Plan-Execute framework that customizes audit strategies based on the unique characteristics of each smart contract. Unlike static, rule-based workflows, our approach iteratively generates and refines audit plans in response to intermediate outputs and newly detected vulnerabilities. To improve reliability, the framework incorporates structured reasoning, prompt optimization, and external tools such as static analyzers and Retrieval-Augmented Generation (RAG). This multi-layered design reduces false positives and enhances the accuracy of vulnerability detection. Experimental results show that SmartAuditFlow achieves 100% accuracy on common vulnerability benchmarks and successfully identifies 13 additional CVEs missed by existing methods. These findings underscore the framework’s adaptability, precision, and practical utility as a robust solution for automated smart contract security auditing.
@@ 1 Introduction
Existing smart contract security auditing primarily relies on manual code reviews by security experts, often supplemented by automated analysis tools Zhang et al. (2023). While valuable, these approaches face significant limitations in the fast-paced blockchain environment. Manual audits, despite their potential for depth, are inherently time-consuming and costly, often creating bottlenecks in development cycles. Furthermore, the continuous evolution of blockchain technology and the emergence of novel attack vectors challenge the ability of human auditors to consistently identify all vulnerabilities Sun et al. (2024a). Expert reviews also rely on individual experience and knowledge, which can introduce subjectivity and gaps in coverage. Conversely, conventional automated tools, while offering speed, frequently suffer from high false positive rates and struggle to detect complex, context-dependent vulnerabilities. Such tools often lack the nuanced understanding of business logic and intricate blockchain-specific security patterns required to identify subtle flaws, such as sophisticated logic errors or vulnerabilities arising from inter-contract interactions, thereby limiting their overall effectiveness.
@@
The advent of Large Language Models (LLMs) presents a promising avenue for advancing smart contract auditing and addressing the aforementioned limitations Liu et al. (2025). LLMs possess the unique ability to process and understand both natural language (e.g., contract specifications, comments) and programming languages (i.e., smart contract code) Liu et al. (2023a). This facilitates a deeper reasoning about the alignment between a contract’s intended functionality and its actual implementation. Their notable capabilities in pattern recognition and contextual analysis are particularly valuable for detecting subtle or novel security vulnerabilities that traditional automated tools might overlook. Furthermore, LLMs are good at reasoning processes and explaining identified security issues in an accessible manner, thereby empowering developers who may not be security specialists to understand and address the findings.
@@
Recent research has shown encouraging results in applying LLMs to smart contract security analysis David et al. (2023); Wei et al. (2023). Nevertheless, substantial challenges hinder the development of fully reliable and effective LLM-based auditing systems. Key issues include: (1) output inconsistency and non-determinism Bender et al. (2021); (2) the difficulty of validating the correctness and completeness of LLM-generated security assessments; and (3) limitations in ensuring comprehensive exploration of potential vulnerabilities. These factors necessitate careful strategies to harness LLM strengths while mitigating their inherent weaknesses. These challenges can be effectively addressed by adopting a workflow-based approach. This methodology decomposes the complex auditing task into manageable, discrete steps, facilitating structured and iterative interactions with LLMs. The efficacy of such structured methods is evident in the improved performance of advanced LLMs in complex reasoning when guided by systematic procedures Zhong et al. (2024); Guan et al. (2024), and is central to agentic workflows like HuggingGPT, Manus, and AutoGPT Shen et al. (2023); Wang et al. (2024), which orchestrate LLMs for sophisticated problem-solving.
@@ 2.1 Smart Contract Auditing
The rapid growth of smart contracts across blockchain networks has brought transformative capabilities, particularly to the Decentralized Finance (DeFi) sector. This rapid adoption, however, has simultaneously highlighted the critical importance of robust security measures, as vulnerabilities within smart contracts have frequently led to substantial financial losses Tolmach et al. (2021); Wei et al. (2023). The DeFi market, for instance, reached a peak total value locked (TVL) of approximately \$179 billion USD on November 9, 2021, with the Ethereum network alone contributing \$108.9 billion USD (around 60% of the total DeFi TVL at the time). As these blockchain-based applications attract significant capital and user engagement, they have also become prime targets for malicious actors. Until the end of 2024, the total amount of money lost by blockchain hackers exceeds 35.32 billion USD, stemming from more than 1800 hacks. A particularly severe recent incident occurred in February 2025, involving the theft of approximately \$1.5 billion USD in Ethereum from the Bybit digital asset exchange, reportedly one of the largest single cryptocurrency heists to date  [^2] https://www.csis.org/analysis/bybit-heist-and-future-us-crypto-regulation.
A smart contract audit is a meticulous and systematic security assessment designed to identify and mitigate vulnerabilities, logical errors, and inefficiencies within the contract’s codebase. Different from foundational blockchain security concerning consensus mechanisms or virtual machines, smart contract security operates at the application layer, which may be more irregular and complex. During an audit, security experts scrutinize all pertinent materials, including source code, whitepapers, and architectural documentation. This process typically involves manual inspection to uncover nuanced flaws that automated tools might miss, complemented by the use of specialized static and dynamic analysis tools to detect common vulnerability patterns and deviations from best practices Ma et al. (2024); Xia et al. (2024). The cost of such audits can range from \$5,000 to over \$50,000 USD, contingent on contract complexity, and demand deep expertise in software engineering, blockchain-specific programming languages (like Solidity), and DeFi protocols. Audit durations can vary from a few days for simple contracts to several weeks or months for complex decentralized applications (dApps).
@@ 2.2 Automated Vulnerability Detection
Traditional automated techniques for smart contract vulnerability detection, including static analysis, symbolic execution, and fuzz testing, have been instrumental in early efforts to secure smart contracts Chaliasos et al. (2024); Durieux et al. (2020). However, these methods face significant challenges related to scalability, computational efficiency, and their ability to comprehensively analyze increasingly complex smart contract systems. A critical study by Zhang et al. Zhang et al. (2023) revealed that 37 widely-used automated detection tools failed to identify over 80% of 516 exploitable vulnerabilities found in 167 real-world smart contracts, underscoring substantial blind spots in conventional approaches. The study also highlighted that the intricacies of detecting complex bugs often necessitate the collaborative efforts of multiple human auditors, further emphasizing the limitations of standalone automated tools and the high cost of thorough manual reviews.
Recent advancements in LLMs offer a promising frontier for enhancing smart contract security analysis David et al. (2023); Wei et al. (2023). LLMs demonstrate strong capabilities in understanding and reasoning about both natural language and programming code. This allows them to potentially bridge the gap between intended contract behavior and actual implementation logic, identifying vulnerabilities, verifying compliance with specifications, and assessing logical correctness with a degree of nuance previously unattainable by purely algorithmic tools. LLMs can excel at reasoning about complex security scenarios, explaining identified vulnerabilities in an understandable manner, evaluating business logic intricacies, and even hypothesizing novel attack vectors Chen et al. (2023a); Sun et al. (2024a). Studies exploring LLMs in the smart contract domain have reported promising results, in some instances demonstrating performance comparable or superior to traditional automated tools. However, the probabilistic nature of LLMs means that their outputs can be inconsistent, and developing robust mechanisms for validating their findings remains an active area of research and a critical consideration for practical deployment Azamfirei et al. (2023); Chen et al. (2023b).
@@ 2.3 Agentic Systems and Workflows
Effectively harnessing LLMs for complex tasks like smart contract auditing requires sophisticated interaction strategies beyond simple querying. Prompt engineering, the art of crafting effective inputs to guide LLM outputs, is a cornerstone of LLM application. It is often more computationally efficient and cost-effective for adapting pre-trained LLMs to specific tasks than full model retraining or extensive fine-tuning, especially for rapid prototyping and iteration. However, LLMs are highly sensitive to prompt phrasing; minor variations in input can lead to significantly different, sometimes degraded, performance. For instance, one study demonstrated accuracy drops of up to 65% across several advanced LLMs due to subtle changes in question framing Mirzadeh et al. (2024). To address this sensitivity and enhance LLM reasoning, researchers have developed advanced prompting methodologies. These include techniques like chain-of-thought (CoT) prompting, which encourages step-by-step reasoning Li et al. (2023); few-shot learning, which provides examples within the prompt Wang and Zhou (2024); and more complex structured reasoning frameworks like Tree of Thought (ToT) Yao et al. (2024) and Graph of Thought (GoT) Besta et al. (2024). These methods aim to elicit more reliable and accurate responses by structuring the LLM’s generation process.
Beyond one attempt advanced prompting, more sophisticated operational paradigms are emerging to tackle multi-step, complex problems. Two prominent approaches are autonomous agents and agentic workflows. Autonomous agents are designed to perceive their environment, make decisions, and take actions dynamically to achieve goals, often involving flexible, self-directed reasoning and tool use Zhuge et al. . They typically require careful design of their action space and decision-making heuristics tailored to their operational context. In contrast, agentic workflows structure problem-solving by decomposing a complex task into a sequence of predefined or dynamically chosen sub-tasks, often involving multiple LLM invocations, potentially with specialized roles or prompts for each step Hong et al. (2024); Zhang et al. (2024). This paradigm emphasizes a more explicit orchestration of LLM capabilities, leveraging human domain expertise in designing the workflow structure and allowing for iterative refinement of both the process and the intermediate results. Agentic workflows can integrate external tools, knowledge bases, and feedback loops, making them well-suited for tasks requiring methodical execution and high reliability, such as security auditing.
@@ 3.2 Plan Phase
The Plan phase is critical for preparing focused and effective tasks for the subsequent Execute Phase. Central to this is guiding the LLM using precisely engineered prompts. Prompts fundamentally shape the LLM’s interpretation of smart contract structures, identification of high-risk functions, and overall focus during analysis. As highlighted by prior research, poorly constructed or misaligned prompts can significantly degrade LLM output quality, potentially leading to incomplete audits or misinterpretation of vulnerabilities Zhong et al. (2024); Guan et al. (2024). Therefore, the systematic generation and selection of optimal prompts for each reasoning step is crucial.
@@ 3.2.1 Initial Analysis and Adaptive Planning Stages (A1-A2)
To achieve A1 and A2 stages, we employ an iterative prompt optimization strategy inspired by recent advancements in automatic prompt engineering Weng et al. (2022); Zhou et al. (2022); Cheng et al. (2023). This strategy aims to discover an optimal instruction $\rho^{*}$ that, when prepended to a given smart contract code $c$, maximizes the quality of the LLM’s output $A$ (e.g., a comprehensive initial analysis or a robust audit plan). Moreover, Stage A1 is enhanced by integrating insights from static analysis tools to ground the LLM’s assessments and improve the accuracy of its initial findings. These insights inform the subsequent prompt generation and planning in later stages.
@@ 3.2.2 Iterative Prompt Optimization Framework
  • Evolution Loop with Mini-Batch Optimization Li et al. (2014); Stapor et al. (2022): Let $\mathcal{U}_{t}$ denote the instruction set at generation $t$. Each iteration of the loop ($t=1\dots T$) proceeds as follows:
@@
This component measures how well the LLM’s generated output $\mathcal{M}(\rho,c)$ aligns with the expected output $A$ from the dataset $\mathcal{D}_{task}$. For smart contract auditing tasks, $A$ might represent an expert-identified set of vulnerabilities, a correctly summarized contract behavior, or a well-structured audit plan component. Alignment can be measured using metrics like BERTScore, ROUGE, or BLEU Vaswani et al. (2017); Es et al. (2024) for textual similarity if $A$ is text, or more domain-specific metrics. We define a domain-specific accuracy metric, drawing upon best practices and guidelines from leading smart contract audit firms (e.g., ConsenSys Diligence, Trail of Bits, or OpenZeppelin).
@@ 3.2.3 Static Analysis Tool Involvement
Widely recognized tools such as Slither, Mythril, Solint, and Oyente automatically detect common vulnerability patterns. However, their utility in our framework extends beyond standalone detection. Specifically, within Stage A1, the outputs from these tools are leveraged to significantly enhance the LLM’s judgements. Among these tools, we selected slither-0.11.0 as our static analysis tool. This choice is motivated by Slither’s active maintenance, frequent updates driven by a large developer community, and its comprehensive suite of nearly 100 distinct vulnerability detectors. In contrast, tools like Mythril primarily focus on low-level EVM bytecode analysis, which, while valuable, offers a different granularity of insight, and other tools such as Solint and Oyente have seen limited updates in recent years. Slither’s proficiency in extracting detailed code syntax and semantic information (e.g., inheritance graphs, function call graphs, state machine representations) has been well-documented in prior research Huang et al. (2023); Ayub et al. (2023).